OpenStudy (anonymous):

The length of a rectangle is five times its width. If the area of the rectangle is 245 ft2 , find its perimeter.

OpenStudy (dayakar):

let w=x ft l = x+5 A = l*w

jhonyy9 (jhonyy9):

@dayakar 5 times not mean 5*x ?

OpenStudy (dayakar):

@jhonyy9, you are right L = 5x